# Reinforcement Learning Teachers of Test Time Scaling

### v1

@@ 1 Introduction
Exploration is one of the critical challenges in reinforcement learning (RL) and has been a core focus of its literature (Bellemare et al., 2016; Salimans and Chen, 2018; Ecoffet et al., 2019). Sparse rewards cannot yield any learning signal unless the agent is already capable of solving the given task at initialization. With the rise of RL for open-ended reasoning (RL reasoning) inducing a new form of language model (LM) scaling beyond prompt-engineering and search (Kojima et al., 2022; Snell et al., 2024), exploration has re-emerged as a key challenge. A canonical motivation for RL is the potential to bootstrap from partial solutions guided by the reward function and learn entirely new tasks from scratch. However, the nature of one-hot correctness rewards used in the RL reasoning framework fails to provide a dense form of guidance, focusing instead on reinforcing correct responses in the initial model’s pool of pass-at-k attempts – without true extrapolation beyond the LM’s initial latent abilities (Yue et al., 2025). As a result, mostly large, already-capable models have been shown to improve consistently beyond cheaper and simpler supervised optimization (Guo et al., 2025).
Due to this fundamental limitation, coupled with RL’s training instability, distillation has emerged as another ubiquitous component of current reasoning paradigms. In this case, the test-time role of LMs trained with RL is to act as a teacher providing instructive reasoning traces for a student to solve new problems. This teacher-student paradigm is widely adopted both to train smaller, less-capable models (Muennighoff et al., 2025; Xu et al., 2025) and even to cold-start future RL iterations for better final convergence with the teacher’s own initial checkpoint acting as the student (Guo et al., 2025; Shen et al., 2025). However, the problem-solving skills reinforced by correctness-based rewards have been shown not to be entirely aligned with the goal of downstream distillation (Li et al., 2025a). To account for this mismatch, current pipelines significantly rely on heuristic-driven postprocessing of the teacher’s outputs for effective student transfer (Guo et al., 2025; Muennighoff et al., 2025; Li et al., 2025a; Labs, 2025).
@@
We will share our code and pretrained checkpoints to facilitate future research in RL reasoning and distillation. In summary, our main contributions are threefold:
@@ 2.1 Reinforcement learning
The RL post-training recipe for inducing reasoning behavior was recently popularized by the DeepSeek R1 line of work (Wang et al., 2023; Shao et al., 2024; Guo et al., 2025). By fine-tuning on a dataset of questions $D=\{q_{1},\dots,q_{N}\}$ with verifiable solutions $\{s_{1},\dots,s_{N}\}$, Guo et al. (2025) show effective “reasoning” behavior emerges out of a 671B-parameter LM (DeepSeek-AI, 2024), significantly pushing its performance on challenging math and coding tasks. Their training is conducted with GRPO (Shao et al., 2024), an online RL algorithm that foregoes the use of a critic model with a simple Monte-Carlo value estimate. GRPO prompts the LM $\pi_{\theta}$ to produce a set of $G>>1$ “grouped” outputs $o_{1},...o_{G}$ for each sampled question $q\in D$, optimizing:
@@
A key component of their design is a system prompt that asks the LM to decompose each generated output $o_{i}$ into two separate formatted sections separated by <think> and <solution> tags, denoted $t_{o_{i}}$ and $s_{o_{i}}$. This structure is forced by assigning rewards $r_{i}=-1$ to unformatted completions, $r_{i}=-0.5$ to wrong but formatted completions, and $r_{i}=1$ only to correct and formatted completions. Training with this strategy, Guo et al. (2025) show the LM’s completion length gradually grows with reflection, verification, and self-correction steps emerging, mirroring human chain-of-thoughts.
@@ 2.2 Supervised distillation
Supervised distillation is another critical step used to train recent reasoning models to complement RL’s shortcomings. For any online RL objective like Equation 1 to avoid collapse, the model must already possess a non-trivial chance of producing correct responses with non-zero gradients at initialization. This defining property makes the RL objective much less applicable than cross-entropy objectives that always include the correct response’s information in the model’s gradients. As a consequence of this dichotomy, distilling the reasoning traces of large RL-trained LMs with supervised learning is not only cheaper but has also been shown to be notably more effective than performing RL itself for inducing reasoning in smaller, less-capable models (Guo et al., 2025; Muennighoff et al., 2025; Ye et al., 2025; Li et al., 2025a; Xu et al., 2025). Furthermore, RL appears prone to instabilities and output degradation, especially during extended training sessions. Due to this second limitation, DeepSeek R1 and several other models (Guo et al., 2025; Shen et al., 2025) perform RL training over multiple iterations. This is done by using the RL-trained models at the end of each intermediate iteration only to, once again, collect distillation datasets used for “cold-starting” their original initial checkpoint and obtain a stronger initialization point for the next RL iteration.
Constructing a dataset of distillation prompts $D_{SD}=\{d_{1},\dots,d_{N}\}$ involves using the RL-trained LM $\pi_{\theta}$ with its reasoning system prompt to answer a corpus of verifiable questions, which can be chosen with several heuristics (Muennighoff et al., 2025; Ye et al., 2025; Li et al., 2025a; Xu et al., 2025). The LM’s output reasoning traces for each question $o\sim\pi_{\theta}(\cdot|q)$ are then filtered by comparing them with the ground-truth solutions to ensure their correctness. Commonly, these reasoning traces are also post-processed via additional “manual” steps of refinements, such as asking other closed-sourced LMs to remove grammatical issues and refactor the reasoning steps into a nicer, consistent format. In fact, Li et al. (2025a) even argues that the structure and format of the thinking data is a critical component to make weaker models actually understand and learn how to reason from distillation, potentially even more important than correctness itself.
@@ 3.4 Pulling everything together: the RLT training paradigm
The RLT framework can be used with any RL algorithm (e.g., (Schulman et al., 2017; Ahmadian et al., 2024)) with minimal modifications to the LM’s conditioning and reward, as described in the above subsections. In this work, we employ the simple GRPO recipe detailed in Section 2, resulting in the following training objective:
@@
where $A^{\mathrm{RLT}}_{i}$ is computed with the normalization strategy defined in Equation 2 using the RLT reward function from Equation 5. Unlike for correctness-based rewards, our learning signal is inherently dense, providing informative rankings to the RLT’s output even before achieving any task expertise. This fundamental difference greatly facilitates our optimization, akin to how heuristically shaped rewards enabled RL agents to learn entirely new behaviors for videogames and robotics tasks (Mnih et al., 2016, 2013).
@@ 4.1 Training, distillation, and evaluation
We train RLTs on the set of questions and solutions selected by Li et al. (2025a) based on their level of challenge. This dataset comprises less than 17K math and coding problems originally used for distilling filtered and post-processed reasoning traces collected from QwQ (Team, 2025a) and DeepSeek R1 (Guo et al., 2025). In contrast, the RLTs we consider are orders-of-magnitude smaller models, all trained starting from the Qwen2.5-7B-Instruct LM (Hui et al., 2024). We precede our RL phase with a short supervised fine-tuning phase to familiarize RLTs with their new system prompt and input format using the open reasoning dataset released by Labs (2025). During RL, we compute the reward for the RLT explanations using another small Qwen-7B model as the student. We train for 250 steps, less than a single epoch, with a small batch size of $256$, a constant learning rate of $1\times 10^{-6}$, and a group size of $64$.
We collect our distillation dataset with the learned RLTs using the same full set of 17K question-solution pairs from training. With the new reasoning traces, we then proceed to fine-tune our students either on this full data or a randomly sampled 1K subset, equating the distillation budget and following the same recipes as our baselines (Muennighoff et al., 2025; Li et al., 2025a). Unlike previous RL distillation pipelines, we do not apply extra postprocessing refinements to improve the quality of the RLT’s reasoning traces, directly using our model’s raw outputs for student fine-tuning. We refer to Appendices A and B for further details regarding our training and distillation phases with complete lists of hyperparameters.
Following prior work (Muennighoff et al., 2025; Labs, 2025; Li et al., 2025a), our main evaluation considers three popular and challenging tasks from the literature: AIME24 (of America, 2024), the set of problems used for the American Invitational Mathematics Examination. MATH 500 (Hendrycks et al., 2021), the set of problems selected by (Lightman et al., 2023) from the canonical competition math benchmark. GPQA Diamond (Rein et al., 2024), the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. We report the completion accuracy of each of our students using Lighteval (Fourrier et al., 2023). When available, we use baseline results reported in prior work, which we found close to our early reproduction attempts. In Appendix C, we extend the experiments in this section by evaluating our models on additional tasks and settings.
@@ 4.2 Test-time reasoning across teachers and students
Our main experiments focus on grounding the effectiveness of RLTs to obtain instructive reasoning traces beyond traditional distillation pipelines. As described in Section 4.1, to construct the student distillation dataset, we use the same starting question-solution pairs as our recent state-of-the-art baselines (Li et al., 2025a; Labs, 2025), with each sample only differing in terms of its reasoning trace. While RLTs could be inexpensively applied to provide explanations of larger corpora, this consistency serves to remove potential confounding factors, other than the quality of the reasoning traces, biasing our experiments and comparisons. For the same reason, we do not retune any hyperparameters for the distillation phase, training students following the same procedure as our baselines based on data size (Li et al., 2025a; Muennighoff et al., 2025).
We compare the RLTs’ explanations with prior approaches, evaluating students fine-tuned on both our full 17K distillation samples and its 1K subset. Our recent baselines all follow a similar recipe of distilling datasets obtained by generating reasoning traces with expensive reasoning models or API calls and postprocessing them with closed-source LMs: s1 (Muennighoff et al., 2025) using traces from Gemini Flash Thinking (Google, 2024), Sky-T1 (Li et al., 2025a) using traces from QwQ (Team, 2025a), and Bespoke (Labs, 2025) using traces from DeepSeek R1 (Guo et al., 2025). Since the Bespoke baseline obtained state-of-the-art results with our same question-solution corpus, we extend its evaluation with new results distilling its processed R1 traces only for our same 1K questions subset, equating its number of datapoints with the other s1 baseline.
@@ 4.3 RLTs to cold-start RL
Our next set of experiments focuses on evaluating the effectiveness of RLTs in providing cold-start data for traditional RL. For this new RL phase, we use our same GRPO implementation with the standard student format and correctness-based rewards described in Section 2. As compared to the RLT hyperparameters, we find a larger batch size of 1024 beneficial to better cope with the increased variance and reward sparsity of traditional RL. We train for a full epoch on the recent RL dataset from Li et al. (2025b) collected by analyzing and selecting an effective subset of the competition math data based on the correlation of individual samples with overall performance improvement.
We compare performing this new RL phase on the Qwen-7B model cold-started from the reasoning traces of our 7B RLT and the postprocessed R1 traces from the Bespoke pipeline, our strongest distillation baseline. Moreover, we also compare a 7B parameters baseline teacher trained with traditional RL as done in prior work (Shen et al., 2025; Guo et al., 2025): effectively performing RL twice on the Qwen model and collecting a dataset at the end of the first iteration to cold-start the second. To construct the cold-starting dataset for this last baseline, we consider either taking the model’s raw output traces, as done with RLTs, or postprocessing them with additional refinements using GPT4.1-mini (Achiam et al., 2023) and following a very similar strategy to the other R1 and QwQ traditional distillation pipelines (Li et al., 2025a; Team, 2025b).
@@
Figure 4: Left: Out-of-distribution performance transferring RLTs to produce new distillation data as compared to students trained on the Li et al. (2025a) corpus and direct RL on the countdown task. Right: Performance after training on different distillation datasets ranked by the RLT reward.
@@ 4.4 Out-of-domain zero-shot transfer
Unlike problem-solving from scratch, we posit that providing effective explanations to given solutions is a much less task-specific skill. Thus, in this subsection, we evaluate how well RLTs can be applied to construct datasets and distill new specialized students in out-of-distribution domains, without any expensive RL retraining. In particular, we focus on the canonical countdown task (Gandhi et al., 2024), asking student models to combine a set of numbers to equal a given target using basic arithmetic operations. We train and test our models on distinct datasets of 16K and 1K automatically-generated question and solution pairs. We compare zero-shot transferring our RLT with transferring the RLT-7B student and the Bespoke-7B baselines from the previous subsections. To ground our results, we also consider performing RL on the countdown task itself (CD RL), training both from the Qwen-7B model and the cold-started Bespoke-7B baseline with the same setup described in Section 4.3.
As shown in the left bar plot of Figure 4, applying RLT distillation zero-shot remarkably achieves even higher performance than direct RL on the countdown task. Interestingly, direct RL appears to provide only marginally better scores than using students distilled from our original set of reasoning questions that do not include any examples of countdown problems (50.8 vs. 49.2). Furthermore, we find there is stark overlap of over 98.5% in the final sets of solved problems between direct RL and the RL-free Bespoke-7B baseline. We find these results in line with prior analysis (Li et al., 2025a; Yue et al., 2025), providing further evidence that the exploration challenge of traditional RL might make most of its benefits come from steering the base model’s distribution toward long-context generation. In contrast, by simplifying the task and foregoing sparse rewards, our RLT appears much more effective – providing countdown-specific traces that allow students to learn new knowledge and solve unseen questions, yielding higher improvements than direct RL even without any teacher training in this new domain.
@@ 4.5 Explanation reward analysis
To analyze the design of the RLT reward function, we start by examining the relationship between the traces’ rewards and the effectiveness of student distillation. In particular, we use our RLT’s checkpoint right before RL training to generate 16 completions for each question-solution pair in our data. We then score all completions with our reward and divide them into groups based on their relative rank for each prompt. Thus, we obtain 16 datasets with different reasoning traces for each question, which we use to train 16 new 7B students from Qwen. As illustrated in the right bar plot of Figure 4, ordering student performance by the respective dataset rank shows a clear correlation between the two, with a Pearson coefficient over 0.89, validating the efficacy of the RLT objective. Additionally, the highest ranked traces of our 7B teacher before any RL remarkably already yield 90% the performance gains of our baseline R1 distillation pipeline (Labs, 2025), showing how even small models already possess latent teaching skills unlocked by our new reward and simplified task formulation.
We also inspect qualitative examples chosen by selecting samples where the reward of the RLT explanations is particularly improved from the baseline R1 distillation pipeline. As shown in Figure 5, we find that the R1 traces with low rewards often try to rely on external tools, such as calculators, and employ language patterns likely idiosyncratic to the training data of the DeepSeek-V3 LM, as sentences with brief humorous comments (DeepSeek-AI, 2024). Instead, the corresponding RLT explanations appear much more grounded and even manage to add new verification steps not considered by R1 to check the final solution. In Appendix 9, we provide additional examples showcasing further qualitative differences of our framework with R1 traces and also specific failure cases from training RLTs without proper balance between each reward component, such as repetitions and overly-long explanations.
Figure 5: Examples comparing the contents from the post-processed R1 traces (Labs, 2025) that were particularly improved by the corresponding RLT explanations as measured by our reward function

### v2

@@ 1 Introduction
Exploration is one of the critical challenges in reinforcement learning (RL) and has been a core focus of its literature [1, 2, 3]. Sparse rewards cannot yield any learning signal unless the agent is already capable of solving the given task at initialization. With the rise of RL for open-ended reasoning (RL reasoning) inducing a new form of language model (LM) scaling beyond prompt-engineering and search [4, 5], exploration has re-emerged as a key challenge. A canonical motivation for RL is the potential to bootstrap from partial solutions guided by the reward function and learn entirely new tasks from scratch. However, the nature of one-hot correctness rewards used in the RL reasoning framework fails to provide a dense form of guidance, focusing instead on reinforcing correct responses in the initial model’s pool of pass-at-k attempts – without true extrapolation beyond the LM’s initial latent abilities [6]. As a result, mostly large, already-capable models have been shown to improve consistently beyond cheaper and simpler supervised optimization [7].
Due to this fundamental limitation, coupled with RL’s training instability, distillation has emerged as another ubiquitous component of current reasoning paradigms. In this case, the test-time role of LMs trained with RL is to act as a teacher providing instructive reasoning traces for a student to solve new problems. This teacher-student paradigm is widely adopted both to train smaller, less-capable models [8, 9] and even to cold-start future RL iterations for better final convergence with the teacher’s own initial checkpoint acting as the student [7, 10]. However, the problem-solving skills reinforced by correctness-based rewards have been shown not to be entirely aligned with the goal of downstream distillation [11]. To account for this mismatch, current pipelines significantly rely on heuristic-driven postprocessing of the teacher’s outputs for effective student transfer [7, 8, 11, 12].
@@
We share our code and pretrained checkpoints [^1] https://github.com/SakanaAI/RLT to facilitate future research in RL reasoning and distillation. In summary, our main contributions are threefold:
@@ 2.1 Reinforcement learning
The RL post-training recipe for inducing reasoning behavior was recently popularized by the DeepSeek R1 line of work [13, 14, 7]. By fine-tuning on a dataset of questions $D=\{q_{1},\dots,q_{N}\}$ with verifiable solutions $\{s_{1},\dots,s_{N}\}$, Guo et al. [7] show effective “reasoning” behavior emerges out of a 671B-parameter LM [15], significantly pushing its performance on challenging math and coding tasks. Their training is conducted with GRPO [14], an online RL algorithm that foregoes the use of a critic model with a simple Monte-Carlo value estimate. GRPO prompts the LM $\pi_{\theta}$ to produce a set of $G>>1$ “grouped” outputs $o_{1},...o_{G}$ for each sampled question $q\in D$, optimizing:
@@
A key component of their design is a system prompt that asks the LM to decompose each generated output $o_{i}$ into two separate formatted sections separated by <think> and <solution> tags, denoted $t_{o_{i}}$ and $s_{o_{i}}$. This structure is forced by assigning rewards $r_{i}=-1$ to unformatted completions, $r_{i}=-0.5$ to wrong but formatted completions, and $r_{i}=1$ only to correct and formatted completions. Training with this strategy, Guo et al. [7] show the LM’s completion length gradually grows with reflection, verification, and self-correction steps emerging, mirroring human chain-of-thoughts.
@@ 2.2 Supervised distillation
Supervised distillation is another critical step used to train recent reasoning models to complement RL’s shortcomings. For any online RL objective like Equation 1 to avoid collapse, the model must already possess a non-trivial chance of producing correct responses with non-zero gradients at initialization. This defining property makes the RL objective much less applicable than cross-entropy objectives that always include the correct response’s information in the model’s gradients. As a consequence of this dichotomy, distilling the reasoning traces of large RL-trained LMs with supervised learning is not only cheaper but has also been shown to be notably more effective than performing RL itself for inducing reasoning in smaller, less-capable models [7, 8, 16, 11, 9]. Furthermore, RL appears prone to instabilities and output degradation, especially during extended training sessions. Due to this second limitation, DeepSeek R1 and several other models [7, 10] perform RL training over multiple iterations. This is done by using the RL-trained models at the end of each intermediate iteration only to, once again, collect distillation datasets used for “cold-starting” their original initial checkpoint and obtain a stronger initialization point for the next RL iteration.
Constructing a dataset of distillation prompts $D_{SD}=\{d_{1},\dots,d_{N}\}$ involves using the RL-trained LM $\pi_{\theta}$ with its reasoning system prompt to answer a corpus of verifiable questions, which can be chosen with several heuristics [8, 16, 11, 9]. The LM’s output reasoning traces for each question $o\sim\pi_{\theta}(\cdot|q)$ are then filtered by comparing them with the ground-truth solutions to ensure their correctness. Commonly, these reasoning traces are also post-processed via additional “manual” steps of refinements, such as asking other closed-sourced LMs to remove grammatical issues and refactor the reasoning steps into a nicer, consistent format. In fact, Li et al. [11] even argues that the structure and format of the thinking data is a critical component to make weaker models actually understand and learn how to reason from distillation, potentially even more important than correctness itself.
@@ 3.4 Pulling everything together: the RLT training paradigm
The RLT framework can be used with any RL algorithm (e.g., [17, 18]) with minimal modifications to the LM’s conditioning and reward, as described in the above subsections. In this work, we employ the simple GRPO recipe detailed in Section 2, resulting in the following training objective:
@@
where $A^{\mathrm{RLT}}_{i}$ is computed with the normalization strategy defined in Equation 2 using the RLT reward function from Equation 5. Unlike for correctness-based rewards, our learning signal is inherently dense, providing informative rankings to the RLT’s output even before achieving any task expertise. This fundamental difference greatly facilitates our optimization, akin to how heuristically shaped rewards enabled RL agents to learn entirely new behaviors for videogames and robotics tasks [19, 20].
@@ 4.1 Training, distillation, and evaluation
We train RLTs on the set of questions and solutions selected by Li et al. [11] based on their level of challenge. This dataset comprises less than 17K math and coding problems originally used for distilling filtered and post-processed reasoning traces collected from QwQ [21] and DeepSeek R1 [7]. In contrast, the RLTs we consider are orders-of-magnitude smaller models, all trained starting from the Qwen2.5-7B-Instruct LM [22]. We precede our RL phase with a short supervised fine-tuning phase to familiarize RLTs with their new system prompt and input format using the open reasoning dataset released by Labs [12]. During RL, we compute the reward for the RLT explanations using another small Qwen-7B model as the student. We train our main models for 125 steps, less than a single epoch, with a batch size of $1024$, a constant learning rate of $1\times 10^{-6}$, and a group size of $64$. We note that we were also able to train RLTs with a smaller batch size of 256 and more steps for faster preliminary experimentation with only slightly inferior results.
We collect our distillation dataset with the learned RLTs using the same full set of 17K question-solution pairs from training. With the new reasoning traces, we then proceed to fine-tune our students either on this full data or a randomly sampled 1K subset, equating the distillation budget and following the same recipes as our baselines [8, 11]. Unlike previous RL distillation pipelines, we do not apply extra postprocessing refinements to improve the quality of the RLT’s reasoning traces, directly using our model’s raw outputs for student fine-tuning. We refer to Appendices A and B for further details regarding our training and distillation phases with complete lists of hyperparameters.
Following prior work [8, 12, 11], our main evaluation considers three popular and challenging tasks from the literature: AIME24 [23], the set of problems used for the American Invitational Mathematics Examination. MATH 500 [24], the set of problems selected by [25] from the canonical competition math benchmark. GPQA Diamond [26], the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. We report the completion accuracy of each of our students using Lighteval [27]. When available, we use baseline results reported in prior work, which we found close to our early reproduction attempts. In Appendix C, we extend the experiments in this section by evaluating our models on additional tasks and settings.
@@ 4.2 Test-time reasoning across teachers and students
Our main experiments focus on grounding the effectiveness of RLTs to obtain instructive reasoning traces beyond traditional distillation pipelines. As described in Section 4.1, to construct the student distillation dataset, we use the same starting question-solution pairs as our recent state-of-the-art baselines [11, 12], with each sample only differing in terms of its reasoning trace. While RLTs could be inexpensively applied to provide explanations of larger corpora, this consistency serves to remove potential confounding factors, other than the quality of the reasoning traces, biasing our experiments and comparisons. For the same reason, we do not retune any hyperparameters for the distillation phase, training students following the same procedure as our baselines based on data size [11, 8].
We compare the RLTs’ explanations with prior approaches, evaluating students fine-tuned on both our full 17K distillation samples and its 1K subset. Our recent baselines all follow a similar recipe of distilling datasets obtained by generating reasoning traces with expensive reasoning models or API calls and postprocessing them with closed-source LMs: s1 [8] using traces from Gemini Flash Thinking [28], Sky-T1 [11] using traces from QwQ [21], and Bespoke [12] using traces from DeepSeek R1 [7]. Since the Bespoke baseline obtained state-of-the-art results with our same question-solution corpus, we extend its evaluation with new results distilling its processed R1 traces only for our same 1K questions subset, equating its number of datapoints with the other s1 baseline.
@@ 4.3 RLTs to cold-start RL
Our next set of experiments focuses on evaluating the effectiveness of RLTs in providing cold-start data for traditional RL. For this new RL phase, we use our same GRPO implementation with the standard student format and correctness-based rewards described in Section 2. As compared to the RLT framework, we find that using a larger batch size of 1024 is significantly more beneficial to better cope with the increased variance and reward sparsity of traditional RL. We train for a full epoch on the recent RL dataset from Li et al. [29] collected by analyzing and selecting an effective subset of the competition math data based on the correlation of individual samples with overall performance improvement.
We compare performing this new RL phase on the Qwen-7B model cold-started from the reasoning traces of our 7B RLT and the postprocessed R1 traces from the Bespoke pipeline, our strongest distillation baseline. Moreover, we also compare a 7B parameters baseline teacher trained with traditional RL as done in prior work [10, 7]: effectively performing RL twice on the Qwen model and collecting a dataset at the end of the first iteration to cold-start the second. To construct the cold-starting dataset for this last baseline, we consider either taking the model’s raw output traces, as done with RLTs, or postprocessing them with additional refinements using GPT4.1-mini [30] and following a very similar strategy to the other R1 and QwQ traditional distillation pipelines [11, 31].
@@
Figure 4: Left: Out-of-distribution performance transferring RLTs to produce new distillation data as compared to students trained on the Li et al. [11] corpus and direct RL on the countdown task. Right: Performance after training on different distillation datasets ranked by the RLT reward.
@@ 4.4 Out-of-domain zero-shot transfer
Unlike problem-solving from scratch, we posit that providing effective explanations to given solutions is a much less task-specific skill. Thus, in this subsection, we evaluate how well RLTs can be applied to construct datasets and distill new specialized students in out-of-distribution domains, without any expensive RL retraining. In particular, we focus on the canonical countdown task [32], asking student models to combine a set of numbers to equal a given target using basic arithmetic operations. We train and test our models on distinct datasets of 16K and 1K automatically-generated question and solution pairs. We compare zero-shot transferring our RLT with transferring the RLT-7B student and the Bespoke-7B baselines from the previous subsections. To ground our results, we also consider performing RL on the countdown task itself (CD RL), training both from the Qwen-7B model and the cold-started Bespoke-7B baseline with the same setup described in Section 4.3.
As shown in the left bar plot of Figure 4, applying RLT distillation zero-shot remarkably achieves even higher performance than direct RL on the countdown task. Interestingly, direct RL appears to provide only marginally better scores than using students distilled from our original set of reasoning questions that do not include any examples of countdown problems (50.8 vs. 49.2). Furthermore, we find there is stark overlap of over 98.5% in the final sets of solved problems between direct RL and the RL-free Bespoke-7B baseline. We find these results in line with prior analysis [11, 6], providing further evidence that the exploration challenge of traditional RL might make most of its benefits come from steering the base model’s distribution toward long-context generation. In contrast, by simplifying the task and foregoing sparse rewards, our RLT appears much more effective – providing countdown-specific traces that allow students to learn new knowledge and solve unseen questions, yielding higher improvements than direct RL even without any teacher training in this new domain.
@@ 4.5 Explanation reward analysis
To analyze the design of the RLT reward function, we start by examining the relationship between the traces’ rewards and the effectiveness of student distillation. In particular, we use our RLT’s checkpoint right before RL training to generate 16 completions for each question-solution pair in our data. We then score all completions with our reward and divide them into groups based on their relative rank for each prompt. Thus, we obtain 16 datasets with different reasoning traces for each question, which we use to train 16 new 7B students from Qwen. As illustrated in the right bar plot of Figure 4, ordering student performance by the respective dataset rank shows a clear correlation between the two, with a Pearson coefficient over 0.89, validating the efficacy of the RLT objective. Additionally, the highest ranked traces of our 7B teacher before any RL remarkably already yield 90% the performance gains of our baseline R1 distillation pipeline [12], showing how even small models already possess latent teaching skills unlocked by our new reward and simplified task formulation.
We also inspect qualitative examples chosen by selecting samples where the reward of the RLT explanations is particularly improved from the baseline R1 distillation pipeline. As shown in Figure 5, we find that the R1 traces with low rewards often try to rely on external tools, such as calculators, and employ language patterns likely idiosyncratic to the training data of the DeepSeek-V3 LM, as sentences with brief humorous comments [15]. Instead, the corresponding RLT explanations appear much more grounded and even manage to add new verification steps not considered by R1 to check the final solution. In Appendix 9, we provide additional examples showcasing further qualitative differences of our framework with R1 traces and also specific failure cases from training RLTs without proper balance between each reward component, such as repetitions and overly-long explanations.
Figure 5: Examples comparing the contents from the post-processed R1 traces [12] that were particularly improved by the corresponding RLT explanations as measured by our reward function